angelaqiu:

A parabola with the equation y=2(x+3)^2-1,determine the equation that will result after each transformation.a compression by a factor of 1/4,followed by a reflection in the y-axis

Vocaloid:

hey there, welcome to QC! it doesn't say whether the compression is horizontal or vertical, so I would guess that it's a vertical compression compressing it by 1/4 would just mean multiply the entire equation by 1/4 and "reflecting across the y-axis" would just be multiplying x (not the entire function, just x) by -1 what do you think the new equation would be after doing these steps? If anything I said was unclear you are welcome to ask for clarification ^_^

Vocaloid:

[actually I just need to make a minor change to what I said before, you don't multiply the entire function by 1/4 just the 2(x+3)^2 part, the -1 stays the same ;__:] sorry about that

angelaqiu:

on the textbook, the answer is y=1/2(x-3)^2-6, but my answer is y=1/2(x-3)^2-1,i'm confused which one is right?

Vocaloid:

huh I have no idea where they would have gotten -6 from

angelaqiu:

Thank you!

Hero:

@angelaqiu, is it possible for you to post a screenshot of the original question?